# Labeling eigenstates of qubit-cavity systems based on the continuity of qubit occupancy: Detecting resonances to higher excited qubit states

### v2

@@ Abstract
We propose a new method for labeling the eigenstates of qubit-cavity systems based on the continuity of the qubit occupancy. The labeled eigenstates give a rough estimation of the evolution of a quantum state under cavity driving. The photon-number dependence of the resonant cavity frequency can be estimated from the labeled eigenenergies, and resonances to higher excited qubit states are visible in the dependence. Our proposed method can be applied to a broader parameter region compared to an existing method. With the proposed method, we investigate the offset charge dependence of the resonances to higher excited states that can induce leakage effects from the computational basis. The results imply that the leakage can occur with only around ten photons.
@@ I Introduction
A system composed of a qubit and a cavity is one of the fundamental setups to study the quantum light-matter interaction, as demonstrated in the seminal Jaynes-Cummings model [1]. Moreover, the system is an essential building block for superconducting quantum computation since the readout of the qubit state can be performed with the setup by the dispersive readout [2, 3, 4]. The dispersive readout is a ubiquitous method in the sense that this readout scheme can be applied to any type of qubit. The importance of qubit-cavity systems is still growing with the development of quantum information technologies.
@@
The qubit used in the superconducting circuits has some complexities that are absent in the Jaynes-Cummings model. The number of states in the qubit is unbounded in general [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. The fluctuation of charged impurities causes the temporal variation of qubit frequency and leads to the decoherence of the qubit [6, 17, 12]. The counter-rotating terms neglected so often [1, 18] can induce the transition to a higher excited qubit state [19]. These complexities should be considered when one studies qubit-cavity systems for practical applications.
@@
Besides, the dispersive readout inherits these complexities. The working principle of the dispersive readout is a qubit-state-dependent frequency shift to the cavity [2, 20, 4, 12]. Higher excited qubit states also affect the frequency shift, and the weight of such excited states depends on the cavity photon number. Consequently, the cavity resonant frequency also depends on the cavity photon number [21, 22, 23]. It is required to evaluate the photon-number dependence of the cavity frequency, especially for operating the readout with high photon numbers. The transition to a higher excited state leads to the leakage of a quantum state from the computational space [19, 24, 25, 23]. Thus, it is also necessary to estimate locations where such transitions occur.
@@
In this paper, we propose a new approach to label the eigenstates of qubit-cavity systems, which can be applied to a broader parameter region compared to an existing method. The photon-number dependence of the cavity frequency can be estimated by labeling the eigenstates with a qubit state and a photon number. A resonance to a higher excited state, which can lead to the transition to the excited state, is visible in the photon-number dependence of the cavity frequency [21, 22].
@@
There already exist some approaches to label the eigenstates [26, 27, 21]. Some of these approaches utilize the conservation of the total number of excitations [26, 27] and cannot be applied to situations where the counter-rotating terms are not negligible. For the method applicable to systems with counter-rotating terms [21], we find instances where the existing method gives suspicious discontinuous photon-number dependence of qubit occupancy. Thus, we develop and propose a new approach that estimates continuous dependence even in such cases. Comparing the photon-number dependence obtained by the existing and the proposed approaches to the evolution under photon driving, we observe that the evolution of the qubit occupancy roughly follows the photon-number dependence given by the proposed method.
@@
With the proposed method, we also investigate the offset charge dependence of resonances to higher qubit excited states in the transmon-cavity system. Contributions from charged impurities are encapsulated in the offset charge. We observe that the locations of the resonances strongly depend on the offset charge and that only around ten photons can induce a transition leading to leakage from the computational space.
@@
The rest of the paper is organized as follows: In Sec. II, we introduce labeling methods used in this paper. The proposed approach is given in this section. In Sec. III, we compare the existing and the proposed labeling methods. The behavior of the proposed method can be controlled by the energy window. We discuss how to choose an appropriate energy window in section IV. In Sec. V, the offset charge dependence of resonances to higher excited qubit states is investigated with the proposed method. The summary is given in Sec. VI.
@@ II Labeling methods
We discuss how to find a labeled state for a qubit state |$\ket{p}_{q}$ and a Fock state |$\ket{n}_{c}$, |$\ket{\overline{p},\overline{n}}$, from the eigenstates of $\hat{H}$. Starting from an initial state |$\ket{\overline{p},\overline{0}}$, the system is expected to follow the labeled states |$\ket{\overline{p},\overline{n}}$ under adiabatic injection of photons into the cavity [26, 22, 21]. The eigenenergy of the labeled eigenstate |$\ket{\overline{p},\overline{n}}$ is denoted by $\varepsilon_{p,n}$. With the labeled eigenenergies, one can regard $(\varepsilon_{p,n+1}-\varepsilon_{p,n})/\hbar$ as the effective cavity frequency when the qubit is initially |$\ket{p}_{q}$ and the cavity photon number is $n$. The photon-number dependence of the cavity frequency is useful information for designing the parameters of the dispersive readout [21, 22]. Besides, a resonance to a higher excited qubit state becomes visible in the photon-number dependence of the cavity frequency.
The simplest approach to find the labeled state |$\ket{\overline{p},\overline{n}}$ is the overlap approach: an eigenstate which has the largest overlap with a product state |$\ket{p}_{q}\otimes\ket{n}_{c}$ is |$\ket{\overline{p},\overline{n}}$. This method does not work in a large $n$ region, as we present in Sec. III. The failure of the overlap approach means that |$\ket{\overline{p},\overline{n}}$ cannot be represented as $\sqrt{1-|\epsilon|^{2}}\ket{p}_{q}\otimes\ket{n}_{c}+\epsilon\ket{\phi}$ for the large $n$ region, where $\epsilon$ is some small constant and |$\ket{\phi}$ is a state orthogonal to |$\ket{p}_{q}\otimes\ket{n}_{c}$.
When the Hamiltonian $\hat{H}$ preserves the total excitation number, i.e., $[\hat{H},\hat{N}_{q}+\hat{c}^{\dagger}\hat{c}]=0$, the method based on block diagonal structure is available [26, 27]. Here, $\hat{N}_{q}$ is the qubit occupation number operator defined by
@@
In such cases, the Hamiltonian has a block diagonal structure with respect to the total number of excitations $\braket{\hat{N}_{q}+\hat{c}^{\dagger}\hat{c}}$. When the eigenstate |$\ket{\lambda}$ belongs to the block corresponds to the total excitation number $M$ and the cavity photon number $\braket{\lambda}{\hat{c}^{\dagger}\hat{c}}{\lambda}$ is close to $n$, |$\ket{\lambda}$ should be |$\ket{\overline{M-n},\overline{n}}$. This method cannot be applied when $\hat{H}$ does not preserve the total excitation number, and the resonance resulting from non-preserving terms has been reported in superconducting qubit systems [19]. Thus, a method applicable to non-preserving systems is preferred.
A recursive approach proposed in Ref. [21] can be applied to non-preserving systems. Starting from |$\ket{\overline{p},\overline{0}}$ which can be obtained by the overlap approach, a state |$\ket{\overline{p},\overline{n}}$ is obtained from |$\ket{\overline{p},\overline{n-1}}$ recursively. In the recursive approach, a state $\hat{c}^{\dagger}\ket{\overline{p},\overline{n-1}}$ is used as a candidate for the next labeled state |$\ket{\overline{p},\overline{n}}$. Therefore, |$\ket{\overline{p},\overline{n}}$ is determined as an eigenstate which has the largest overlap with the state $\hat{c}^{\dagger}\ket{\overline{p},\overline{n-1}}$. This labeling method is used to analyze the readout dynamics of the transmon qubit with high-power input light, and the effective cavity frequency obtained by this method reflects the resonance coming from non-preserving terms [21, 22].
The recursive approach has shown good performance in the previous studies [21, 22]. However, we find instances where the recursive approach gives the discontinuous dependence of physical quantities on photon number $n$. Consequently, we propose an approach that estimates the continuous dependence of physical quantities even in the case where the recursive approach does not.
@@ III Comparison of labeling methods
In this section, we compare three labeling methods presented in Sec. II, i.e., the overlap, the recursive [21], and the proposed methods. In the comparison, we choose the transmon as the qubit whose Hamiltonian is given by [5, 10, 12, 11]
@@
Figure 1: Photon-number dependence of the cavity frequencies in the transmon-cavity system obtained by the three labeling methods. The recursive calculations start from |$\ket{\overline{g},\overline{0}}$. For the system parameters, we use $(E_{C}/\hbar\omega_{c},E_{J}/\hbar\omega_{c},g/\omega_{c},N_{g})=(5.0\times 10^{-2},1.6,2.5\times 10^{-2},0.0)$. The energy window $\delta$ is set to $1.0\times 10^{-2}\hbar\omega_{c}$ for the proposed method.
Figure 1 gives the photon-number dependence of the cavity frequencies obtained by the three labeling methods. The recursive calculations start from |$\ket{\overline{g},\overline{0}}$. We use $(E_{C}/\hbar\omega_{c},E_{J}/\hbar\omega_{c},g/\omega_{c},N_{g})=(5.0\times 10^{-2},1.6,2.5\times 10^{-2},0.0)$ for the system parameters. The energy window is set to $\delta=1.0\times 10^{-2}\hbar\omega_{c}$. The three labeling methods result in the same labeling up to $n\sim 150$. The sharp peak around $n\sim 50$ comes from the resonance with the higher excited qubit states. The transition to higher excited state occurs when the cavity photon number stays near the resonant peak. However, such transition can be avoided by quickly passing the resonant point [21, 22]. Around $n\sim 150$, the frequency given by the overlap method gives a discontinuous jump. After the discontinuous jump, the cavity frequency obtained by the overlap method shows non-smooth dependence. The observation of unphysical non-smooth behavior means that the overlap method fails to trace the labeled states |$\ket{\overline{g},\overline{n}}$, and the failure implies that the labeled states |$\ket{\overline{g},\overline{n}}$ are not close to |$\ket{g}_{q}\otimes\ket{n}_{c}$ for $n\gtrsim 150$.
The other two methods give different cavity frequencies after $n\sim 180$. The cavity frequency given by the recursive method shows a sudden drop around $n\sim 180$, while that estimated by the proposed method indicates the presence of the resonance at the point. Since these two behaviors are possible, we cannot exclude either of the behaviors only with the data. Consequently, we choose the better labeling method by comparing the qubit occupation numbers $\braket{\hat{N}_{q}}$ of the labeled states and that evolved under cavity driving.
@@
where $\braket{\hat{A}(t)}_{U}$ stands for $\braket{\psi(t)}{\hat{U}^{\dagger}(t)\hat{A}\hat{U}(t)}{\psi(t)}$ [22]. For the numerical simulation of the dynamics in this section, we use the Fock state of the cavity up to |$\ket{150}_{c}$, and the drive frequency is set to $\omega_{d}/\omega_{c}=1.0015$.
Figure 2: Photon-number dependence of the qubit occupation number in the transmon-cavity system obtained by the two labeling methods (recursive and proposed) and the evolution under cavity driving. The recursive calculations start from |$\ket{\overline{g},\overline{0}}$. An initial state for the dynamics is also set to |$\ket{\overline{g},\overline{0}}$. The parameters of the system are the same as those in Fig. 1. The amplitude of the driving field $E$ is set to $5.0\times 10^{-3}\omega_{c}$, and the drive frequency $\omega_{d}$ is set to $1.0015\omega_{c}$.
@@
Under cavity driving, the evolution of the qubit occupation number roughly follows the photon-number dependence obtained by the proposed method apart from the resonant peaks. Therefore, we conclude that the proposed method can be applied to a broader parameter region compared to the recursive method. We present another numerical example with parameters $(E_{C}/\hbar\omega_{c},E_{J}/\hbar\omega_{c},g/\omega_{c},N_{g},\delta/\hbar\omega_{c},E/\omega_{c},\omega_{d}/\omega_{c})=(5.0\times 10^{-2},2.0,1.0\times 10^{-2},0.1,1.0\times 10^{-2},5.0\times 10^{-3},1.0002)$ starting from the excited state |$\ket{\bar{e},\bar{0}}$. For this case with a finite charge offset, the same conclusion is obtained as summarized in Fig. 3. The recursive method fails to follow the evolution of states at the resonant point, while the proposed method succeeds. The proposed method can access the effective cavity frequency beyond the resonant peak.
@@
Figure 3: Photon-number dependence of (a) the cavity frequency and (b) the qubit occupation number in the transmon-cavity system obtained by the two labeling methods. For the qubit occupation number, the evolution under cavity driving is also plotted. The recursive calculations start from |$\ket{\overline{e},\overline{0}}$. An initial state for the dynamics is also set to |$\ket{\overline{e},\overline{0}}$. For the calculation, we use $(E_{C}/\hbar\omega_{c},E_{J}/\hbar\omega_{c},g/\omega_{c},N_{g},\delta/\hbar\omega_{c},E/\omega_{c},\omega_{d}/\omega_{c})=(5.0\times 10^{-2},2.0,1.0\times 10^{-2},0.1,1.0\times 10^{-2},5.0\times 10^{-3},1.0002)$.
IV How to choose appropriate energy window
@@
Figure 4: Photon-number dependence of (a) the cavity frequency and (b) the qubit occupation number in the transmon-cavity system obtained by the proposed labeling methods with different energy windows $\delta$. For the qubit occupation number, the evolution under cavity driving is also plotted. The labeling method starts from |$\ket{\overline{g},\overline{0}}$. An initial state for the dynamics is also set to |$\ket{\overline{g},\overline{0}}$. For the calculation, we use $(E_{C}/\hbar\omega_{c},E_{J}/\hbar\omega_{c},g/\omega_{c},N_{g},E/\omega_{c},\omega_{d}/\omega_{c})=(5.0\times 10^{-2},1.6,2.5\times 10^{-2},0.1,1.0\times 10^{-2},1.001)$.
@@
Compared to the recursive approach, our proposed method introduces an additional parameter, the energy window $\delta$. This parameter should be large enough to include the eigenenergy affected by a higher excited state within the energy window and small enough to exclude eigenstates whose qubit occupancies are accidentally close to the target value. The additional parameter can adjust the behavior of the labeling method, and thus the proposed method acquires adaptability. In this section, we discuss how to choose an appropriate energy window $\delta$.
@@
To discuss the appropriate choice of the energy window, we observe the photon-number dependence of the effective cavity frequency and the qubit occupancy given by the proposed labeling method with different energy windows, as shown in Fig. 4. We use the system parameters $(E_{C}/\hbar\omega_{c},E_{J}/\hbar\omega_{c},g/\omega_{c},N_{g})=(5.0\times 10^{-2},1.6,2.5\times 10^{-2},0.1)$ and three distinct energy windows $1.0\times 10^{-2}\hbar\omega_{c}$, $1.5\times 10^{-2}\hbar\omega_{c}$, and $2.0\times 10^{-2}\hbar\omega_{c}$. For the largest energy window $2.0\times 10^{-2}\hbar\omega_{c}$, even the effective cavity frequency shows a discontinuous jump, and this jump suggests that the selected energy window is too large and includes an unnecessary state. For the other two energy windows, a discontinuous behavior is not visible in the cavity frequency and the qubit occupancy except for resonant peaks.

### v3

@@ Abstract
We propose a new method for labeling the eigenstates of qubit-cavity systems based on the continuity of the qubit occupancy. The labeled eigenstates give a rough estimate of the evolution of a quantum state under cavity driving. The photon-number dependence of the resonant cavity frequency can be estimated from the labeled eigenenergies, and resonances to higher excited qubit states are visible in the dependence. Our proposed method can be applied to a broader situation compared to an existing method. With the proposed method, we investigate the offset-charge dependence of the resonances to higher excited states that can induce leakage effects from the computational basis. The results imply that the leakage can occur with only around ten photons.
@@ I Introduction
A system composed of a qubit and a cavity is one of the fundamental setups to study the quantum light-matter interaction, as demonstrated in the seminal Jaynes-Cummings model Jaynes and Cummings (1963). Moreover, the system is an essential building block for superconducting quantum computation since the readout of the qubit state can be performed with the setup by the dispersive readout Blais et al. (2004); Wallraff et al. (2004); Vijay et al. (2011). The dispersive readout is a ubiquitous method in the sense that this readout scheme can be applied to any type of qubit. The importance of qubit-cavity systems is still growing with the development of quantum information technologies.
@@
The qubit used in the superconducting circuits has some complexities that are absent in the Jaynes-Cummings model. The number of states in the qubit is unbounded in general Bouchiat et al. (1998); Nakamura et al. (1999); Orlando et al. (1999); Yan et al. (2016); Martinis et al. (2002); Koch et al. (2007); Vool and Devoret (2017); Blais et al. (2021); Manucharyan et al. (2009); Kalashnikov et al. (2020); Patel et al. (2024); Brosco et al. (2024). The fluctuation of charged impurities causes the temporal variation of qubit frequency and leads to the decoherence of the qubit Nakamura et al. (1999); Metcalfe et al. (2007); Blais et al. (2021). The counter-rotating terms neglected so often Jaynes and Cummings (1963); Tavis and Cummings (1968) can induce the transition to a higher excited qubit state Sank et al. (2016). These complexities should be considered when one studies qubit-cavity systems for practical applications.
@@
Besides, the dispersive readout inherits these complexities. The working principle of the dispersive readout is a qubit-state-dependent frequency shift to the cavity Blais et al. (2004, 2007); Vijay et al. (2011); Blais et al. (2021). Higher excited qubit states also affect the frequency shift, and the weight of such excited states depends on the cavity photon number. Consequently, the cavity resonant frequency also depends on the cavity photon number Shillito et al. (2022); Goto and Koshino (2023); Khezri et al. (2023). It is required to evaluate the photon-number dependence of the cavity frequency, especially for operating the readout with high photon numbers. The transition to a higher excited state leads to the leakage of a quantum state from the computational space Sank et al. (2016); Lescanne et al. (2019); Verney et al. (2019); Khezri et al. (2023). Thus, it is also necessary to estimate locations where such transitions occur.
@@
In this paper, we propose a new approach to label the eigenstates of qubit-cavity systems, which can be applied to a broader situation compared to an existing method. The photon-number dependence of the cavity frequency can be estimated by labeling the eigenstates with a qubit state and a photon number. A resonance to a higher excited state, which can lead to the transition to the excited state, is visible in the photon-number dependence of the cavity frequency Shillito et al. (2022); Goto and Koshino (2023).
@@
There already exist some approaches to label the eigenstates Boissonneault et al. (2010); Khezri et al. (2016); Shillito et al. (2022); Xiao et al. (2023). Floquet modes in driven systems are labeled Xiao et al. (2023) as well as the eigenstates of undriven systems. Some of these approaches utilize the conservation of the total number of excitations Boissonneault et al. (2010); Khezri et al. (2016) and cannot be applied to situations where the counter-rotating terms are not negligible. For the method applicable to systems with counter-rotating terms Shillito et al. (2022), we find instances where the existing method gives discontinuous photon-number dependence of qubit occupancy. In such instances, the obtained labeling follows a resonant transition, and one cannot estimate the cavity frequency for a case where the transition is avoided by quickly passing the resonant point. Thus, we develop and propose a new approach that estimates continuous dependence even in such cases. Comparing the photon-number dependence obtained by the existing and the proposed approaches to the evolution under strong photon driving enough to avoid the resonant transition, we observe that the evolution of the qubit occupancy roughly follows the photon-number dependence given by the proposed method.
@@
With the proposed method, we also investigate the offset-charge dependence of resonances to higher qubit excited states in the transmon-cavity system. Contributions from charged impurities are encapsulated in the offset charge. We observe that the locations of the resonances strongly depend on the offset charge and that only around ten photons can induce a transition leading to leakage from the computational space.
@@
The rest of the paper is organized as follows: In Sect. II, we introduce labeling methods used in this paper. The proposed approach is given in this section. In Sect. III, we compare the existing and the proposed labeling methods. The behavior of the proposed method can be controlled by the energy window. We discuss how the choice of the energy window affects the labeling in Sect. IV. In Sect. V, the offset-charge dependence of resonances to higher excited qubit states is investigated with the proposed method. The summary is given in Sect. VI.
@@ II Labeling methods
We discuss how to find a labeled state for a qubit state |$\ket{p}_{q}$ and a Fock state |$\ket{n}_{c}$, |$\ket{\overline{p},\overline{n}}$, from the eigenstates of $\hat{H}$. Starting from an initial state |$\ket{\overline{p},\overline{0}}$, the system is expected to follow the labeled states |$\ket{\overline{p},\overline{n}}$ under adiabatic injection of photons into the cavity Boissonneault et al. (2010); Goto and Koshino (2023); Shillito et al. (2022). The eigenenergy of the labeled eigenstate |$\ket{\overline{p},\overline{n}}$ is denoted by $\varepsilon_{p,n}$. With the labeled eigenenergies, one can regard $(\varepsilon_{p,n+1}-\varepsilon_{p,n})/\hbar$ as the effective cavity frequency when the qubit is initially |$\ket{p}_{q}$ and the cavity photon number is $n$. The photon-number dependence of the cavity frequency is useful information for designing the parameters of the dispersive readout Shillito et al. (2022); Goto and Koshino (2023). Besides, a resonance to a higher excited qubit state becomes visible in the photon-number dependence of the cavity frequency.
The simplest approach to find the labeled state |$\ket{\overline{p},\overline{n}}$ is the overlap approach: an eigenstate which has the largest overlap with a product state |$\ket{p}_{q}\otimes\ket{n}_{c}$ is |$\ket{\overline{p},\overline{n}}$. This method does not work in a large $n$ region, as we will present in Sect. III. The failure of the overlap approach means that |$\ket{\overline{p},\overline{n}}$ cannot be represented as $\sqrt{1-|\epsilon|^{2}}\ket{p}_{q}\otimes\ket{n}_{c}+\epsilon\ket{\phi}$ for the large $n$ region, where $\epsilon$ is some small constant and |$\ket{\phi}$ is a state orthogonal to |$\ket{p}_{q}\otimes\ket{n}_{c}$.
When the Hamiltonian $\hat{H}$ preserves the total excitation number, i.e., $[\hat{H},\hat{N}_{q}+\hat{c}^{\dagger}\hat{c}]=0$, the method based on block diagonal structure is available Boissonneault et al. (2010); Khezri et al. (2016). Here, $\hat{N}_{q}$ is the qubit occupation number operator defined by
@@
In such cases, the Hamiltonian has a block diagonal structure with respect to the total number of excitations $\braket{\hat{N}_{q}+\hat{c}^{\dagger}\hat{c}}$. When the eigenstate |$\ket{\lambda}$ belongs to the block corresponds to the total excitation number $M$ and the cavity photon number $\braket{\lambda}{\hat{c}^{\dagger}\hat{c}}{\lambda}$ is close to $n$, |$\ket{\lambda}$ should be |$\ket{\overline{M-n},\overline{n}}$. This method cannot be applied when $\hat{H}$ does not preserve the total excitation number, and the resonance resulting from non-preserving terms has been reported in superconducting qubit systems Sank et al. (2016). Thus, a method applicable to non-preserving systems is preferred.
A recursive approach proposed in Ref. Shillito et al. (2022) can be applied to non-preserving systems. Starting from |$\ket{\overline{p},\overline{0}}$ which can be obtained by the overlap approach, a state |$\ket{\overline{p},\overline{n}}$ is obtained from |$\ket{\overline{p},\overline{n-1}}$ recursively. In the recursive approach, a state $\hat{c}^{\dagger}\ket{\overline{p},\overline{n-1}}$ is used as a candidate for the next labeled state |$\ket{\overline{p},\overline{n}}$. Therefore, |$\ket{\overline{p},\overline{n}}$ is determined as an eigenstate which has the largest overlap with the state $\hat{c}^{\dagger}\ket{\overline{p},\overline{n-1}}$. This labeling method is used to analyze the readout dynamics of the transmon qubit with high-power input light, and the effective cavity frequency obtained by this method reflects the resonance coming from non-preserving terms Shillito et al. (2022); Goto and Koshino (2023).
The recursive approach has shown good performance in the previous studies Shillito et al. (2022); Goto and Koshino (2023). However, we find instances where the recursive approach gives the discontinuous dependence of physical quantities on photon number $n$. Consequently, we propose an approach that estimates the continuous dependence of physical quantities even in the case where the recursive approach does not.
@@ III Comparison of labeling methods
In this section, we compare three labeling methods presented in Sect. II, i.e., the overlap, the recursive Shillito et al. (2022), and the proposed methods. In the comparison, we choose the transmon as the qubit whose Hamiltonian is given by Bouchiat et al. (1998); Koch et al. (2007); Blais et al. (2021); Vool and Devoret (2017)
@@
Figure 1: (Color online) Photon-number dependence of the cavity frequencies in the transmon-cavity system obtained by the three labeling methods. The recursive calculations start from |$\ket{\overline{g},\overline{0}}$. For the system parameters, we use $(E_{C}/\hbar\omega_{c},E_{J}/\hbar\omega_{c},g/\omega_{c},N_{g})=(5.0\times 10^{-2},1.6,2.5\times 10^{-2},0.0)$. The energy window $\delta$ is set to $1.0\times 10^{-2}\hbar\omega_{c}$ for the proposed method.
Figure 1 gives the photon-number dependence of the cavity frequencies obtained by the three labeling methods. The recursive calculations start from |$\ket{\overline{g},\overline{0}}$. We use $(E_{C}/\hbar\omega_{c},E_{J}/\hbar\omega_{c},g/\omega_{c},N_{g})=(5.0\times 10^{-2},1.6,2.5\times 10^{-2},0.0)$ for the system parameters. With the parameters, the occupations of |$\ket{10}_{t}$ and |$\ket{-10}_{t}$ in the qubit eigenstate |$\ket{10}_{q}$ are only on the order of $10^{-10}$. The qubit eigenstates energetically lower than |$\ket{10}_{q}$ are little affected by the truncation of the charge basis. The energy window is set to $\delta=1.0\times 10^{-2}\hbar\omega_{c}$. The three labeling methods result in the same labeling up to $n\sim 150$. The sharp peak around $n\sim 50$ comes from the resonance with the higher excited qubit states. The transition to higher excited state occurs when the cavity photon number stays near the resonant peak. However, such transition can be avoided by quickly passing the resonant point Shillito et al. (2022); Goto and Koshino (2023). Around $n\sim 150$, the frequency given by the overlap method gives a discontinuous jump. After the discontinuous jump, the cavity frequency obtained by the overlap method shows non-smooth dependence. The observation of unphysical non-smooth behavior means that the overlap method fails to trace the labeled states |$\ket{\overline{g},\overline{n}}$, and the failure implies that the labeled states |$\ket{\overline{g},\overline{n}}$ are not close to |$\ket{g}_{q}\otimes\ket{n}_{c}$ for $n\gtrsim 150$.
The other two methods give different cavity frequencies after $n\sim 180$. The cavity frequency given by the recursive method shows a sudden drop around $n\sim 180$, while that estimated by the proposed method indicates the presence of the resonance at the point. This observation suggests that the recursive method follows a resonant transition, while the proposed method passes through the transition. Next, we try to confirm that the proposed method can actually avoid the resonant transitions by comparing the qubit occupation numbers $\braket{\hat{N}_{q}}$ of the labeled states and that evolved under cavity driving.
@@
where $\braket{\hat{A}(t)}_{U}$ stands for $\braket{\psi(t)}{\hat{U}^{\dagger}(t)\hat{A}\hat{U}(t)}{\psi(t)}$ Goto and Koshino (2023). For the numerical simulation of the dynamics in this section, we use the Fock state of the cavity up to |$\ket{150}_{c}$, and the drive frequency is set to $\omega_{d}/\omega_{c}=1.0015$.
Figure 2: (Color online) Photon-number dependence of the qubit occupation number in the transmon-cavity system obtained by the two labeling methods (recursive and proposed) and the evolution under cavity driving. The recursive calculations start from |$\ket{\overline{g},\overline{0}}$. An initial state for the dynamics is also set to |$\ket{\overline{g},\overline{0}}$. The parameters of the system are the same as those in Fig. 1. The amplitude of the driving field $E$ is set to $5.0\times 10^{-3}\omega_{c}$, and the drive frequency $\omega_{d}$ is set to $1.0015\omega_{c}$.
@@
Under cavity driving, the evolution of the qubit occupation number roughly follows the photon-number dependence obtained by the proposed method apart from the resonant peaks. Therefore, we conclude that the proposed method can pass through the resonant transition and be applied to a broader situation compared to the recursive method. We present another numerical example with parameters $(E_{C}/\hbar\omega_{c},E_{J}/\hbar\omega_{c},g/\omega_{c},N_{g},\delta/\hbar\omega_{c},E/\omega_{c},\omega_{d}/\omega_{c})=(5.0\times 10^{-2},2.0,1.0\times 10^{-2},0.1,1.0\times 10^{-2},5.0\times 10^{-3},1.0002)$ starting from the excited state |$\ket{\bar{e},\bar{0}}$. For this case with a finite charge offset, the same conclusion is obtained as summarized in Fig. 3. The recursive method follows a resonant transition, while the proposed method avoids the transition. The proposed method can access the effective cavity frequency when the system passes through the resonant transition by sufficiently strong driving.
@@
Figure 3: (Color online) Photon-number dependence of (a) the cavity frequency and (b) the qubit occupation number in the transmon-cavity system obtained by the two labeling methods. For the qubit occupation number, the evolution under cavity driving is also plotted. The recursive calculations start from |$\ket{\overline{e},\overline{0}}$. An initial state for the dynamics is also set to |$\ket{\overline{e},\overline{0}}$. For the calculation, we use $(E_{C}/\hbar\omega_{c},E_{J}/\hbar\omega_{c},g/\omega_{c},N_{g},\delta/\hbar\omega_{c},E/\omega_{c},\omega_{d}/\omega_{c})=(5.0\times 10^{-2},2.0,1.0\times 10^{-2},0.1,1.0\times 10^{-2},5.0\times 10^{-3},1.0002)$.
IV energy-window dependence
@@
Figure 4: (Color online) Photon-number dependence of (a) the cavity frequency and (b) the qubit occupation number in the transmon-cavity system obtained by the proposed labeling methods with different energy windows $\delta$. For the qubit occupation number, the evolution under cavity driving is also plotted. The results obtained by the recursive labeling method are plotted for comparison. The labeling method starts from |$\ket{\overline{g},\overline{0}}$. An initial state for the dynamics is also set to |$\ket{\overline{g},\overline{0}}$. For the calculation, we use $(E_{C}/\hbar\omega_{c},E_{J}/\hbar\omega_{c},g/\omega_{c},N_{g},E/\omega_{c},\omega_{d}/\omega_{c})=(5.0\times 10^{-2},1.6,2.5\times 10^{-2},0.1,1.0\times 10^{-2},1.001)$.
@@
Compared to the recursive approach, our proposed method introduces an additional parameter, the energy window $\delta$. This parameter should be large enough to include the eigenenergy affected by a higher excited state within the energy window and small enough to exclude eigenstates whose qubit occupancies are accidentally close to the target value. The additional parameter can adjust the behavior of the labeling method, and thus the proposed method acquires adaptability. In this section, we discuss how the energy window $\delta$ affects the labeling.
@@
To discuss the effects of the energy window, we observe the photon-number dependence of the effective cavity frequency and the qubit occupancy given by the proposed labeling method with different energy windows, as shown in Fig. 4. We use the system parameters $(E_{C}/\hbar\omega_{c},E_{J}/\hbar\omega_{c},g/\omega_{c},N_{g})=(5.0\times 10^{-2},1.6,2.5\times 10^{-2},0.1)$ and three distinct energy windows $1.0\times 10^{-2}\hbar\omega_{c}$, $1.5\times 10^{-2}\hbar\omega_{c}$, and $2.0\times 10^{-2}\hbar\omega_{c}$. For the largest energy window $2.0\times 10^{-2}\hbar\omega_{c}$, the effective cavity frequency shows a discontinuous jump, and this jump suggests that the selected energy window is so large that a highly excited qubit state is included in the energy window. The labeling with this energy window is very close to the labeling by the recursive method. The recursive method can detect the resonance around $n\sim 40$, while it cannot detect those around $n\sim 110$ and $n\sim 130$. For the other two energy windows, a discontinuous behavior is not visible in the cavity frequency and the qubit occupancy except for resonant peaks.